# Basis Selection: Low-Rank Decomposition of Pretrained Large Language Models for Target Applications

### v2

@@ 5 Conclusion
The significant size of large language models leads to high inference costs and demands substantial computing resources. To mitigate these issues, we focus on compressing large language models to meet the specific requirements of target applications. Our approach involves examining these models through the lens of matrix factorization. By viewing the weight matrix of large language models as a linear combination of a group of bases, we have identified that pretrained models often contain many redundant bases that are less useful for target applications. To address this, we propose Basel, a compression algorithm that evaluates the importance of each base for target applications and prunes those that are less significant. Experimental results demonstrate that Basel significantly outperforms state-of-the-art low-rank compression algorithms in achieving deep compression. Basel greatly reduces the inference cost of large language models, making them more accessible and practical for a wider range of applications. This advancement has the potential to democratize the use of large language models, facilitating their adoption and integration across diverse fields and industries.
Список литературы
  • Acharya et al. (2019) Anish Acharya, Rahul Goel, Angeliki Metallinou, and Inderjit Dhillon. Online Embedding Compression for Text Classification Using Low Rank Matrix Factorization. In AAAI, 2019.
  • An et al. (2024) Yongqi An, Xu Zhao, Tao Yu, Ming Tang, and Jinqiao Wang. Fluctuation-based Adaptive Structured Pruning for Large Language Models. In AAAI, 2024.
  • Austin et al. (2021) Jacob Austin, Augustus Odena, Maxwell Nye, Maarten Bosma, Henryk Michalewski, David Dohan, Ellen Jiang, Carrie Cai, Michael Terry, Quoc Le, and Charles Sutton. Program Synthesis with Large Language Models. arXiv preprint arXiv:2108.07732, 2021.
  • Chen et al. (2021a) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, et al. Evaluating Large Language Models Trained on Code. arXiv preprint arXiv:2107.03374, 2021a.
  • Chen et al. (2018) Patrick Chen, Si Si, Yang Li, Ciprian Chelba, and Cho-Jui Hsieh. GroupReduce: Block-Wise Low-Rank Approximation for Neural Language Model Shrinking. In NeurIPS, 2018.
  • Chen et al. (2021b) Patrick Chen, Hsiang-Fu Yu, Inderjit Dhillon, and Cho-Jui Hsieh. DRONE: Data-Aware Low-Rank Compression for Large NLP Models. In NeurIPS, 2021b.
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. Training Verifiers to Solve Math Word Problems. arXiv preprint arXiv:2110.14168, 2021.
  • Denton et al. (2014) Emily L Denton, Wojciech Zaremba, Joan Bruna, Yann LeCun, and Rob Fergus. Exploiting Linear Structure within Convolutional Networks for Efficient Evaluation. NeurIPS, 2014.
  • Dettmers et al. (2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer. QLoRA: Efficient Finetuning of Quantized LLMs. In NeurIPS, 2023.
  • Golub & Van Loan (1996) Gene H. Golub and Charles F. Van Loan. Matrix Computations. Johns Hopkins University Press, Baltimore and London, 1996.
  • Hendrycks et al. (2021) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Song, and Jacob Steinhardt. Measuring Mathematical Problem Solving with the MATH Dataset. NeurIPS, 2021.
  • Hsu et al. (2022) Yen-Chang Hsu, Ting Hua, Sungen Chang, Qian Lou, Yilin Shen, and Hongxia Jin. Language Model Compression with Weighted Low-Rank Factorization. In ICLR, 2022.
  • Jaderberg et al. (2014) Max Jaderberg, Andrea Vedaldi, and Andrew Zisserman. Speeding Up Convolutional Neural Networks with Low Rank Expansions. arXiv preprint arXiv:1405.3866, 2014.
  • Noach & Goldberg (2020) Matan Ben Noach and Yoav Goldberg. Compressing pre-trained language models by matrix decomposition. In AACL/IJCNLP, 2020.
  • Povey et al. (2018) Daniel Povey, Gaofeng Cheng, Yiming Wang, Ke Li, Hainan Xu, Mahsa Yarmohammadi, and Sanjeev Khudanpur. Semi-Orthogonal Low-Rank Matrix Factorization for Deep Neural Networks. In INTERSPEECH, 2018.
  • Sharma et al. (2024) Pratyusha Sharma, Jordan T. Ash, and Dipendra Misra. The Truth is in There: Improving Reasoning in Language Models with Layer-Selective Rank Reduction. In ICLR, 2024.
  • Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. Llama: Open and Efficient Foundation Language Models. arXiv preprint arXiv:2302.13971, 2023a.
  • Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. Llama 2: Open Foundation and Finetuned Chat Models. arXiv preprint arXiv:2307.09288, 2023b.
  • Xue et al. (2013) Jian Xue, Jinyu Li, and Yifan Gong. Restructuring of Deep Neural Network Acoustic Models with Singular Value Decomposition. In INTERSPEECH, 2013.
  • Yu & Wu (2023) Hao Yu and Jianxin Wu. Compressing Transformers: Features Are Low-Rank, But Weights Are Not! In AAAI, 2023.
  • Zhang et al. (2015) Xiangyu Zhang, Jianhua Zou, Kaiming He, and Jian Sun. Accelerating very deep convolutional networks for classification and detection. IEEE Transactions on Pattern Analysis and Machine Intelligence, 38(10):1943–1955, 2015.

### v4

@@ 5 Conclusion
The significant size of large language models leads to high inference costs and demands substantial computing resources. To mitigate these issues, we focus on compressing large language models to meet the specific requirements of target applications. Our approach involves examining these models through the lens of matrix factorization. By viewing the weight matrix of large language models as a linear combination of a group of bases, we have identified that pretrained models often contain many redundant bases that are less useful for target applications. To address this, we propose Basel, a compression algorithm that evaluates the importance of each base for target applications and prunes those that are less significant. Experimental results demonstrate that Basel significantly outperforms baseline low-rank compression algorithms in achieving deep compression. Basel greatly reduces the inference cost of large language models, making them more accessible and practical for a wider range of applications. This advancement has the potential to democratize the use of large language models, facilitating their adoption and integration across diverse fields and industries.
Acknowledgments
This research is partially supported by Meta and by a faculty startup grant from Iowa State University. Computational resources are provided by Meta and HPC@ISU. The resources from HPC@ISU include equipment funded by the U.S. National Science Foundation under MRI Grant Nos. 1726447 and 2018594.
Appendix
Таблица 4: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms on the mathematical reasoning task.
SVD Model Size (billion) 6.74 5.02 3.18 1.73 1.11 0.56
GSM8K Acc (%) 66.4 63.0 61.0 53.9 32.9 11.9
MATH Acc (%) 20.6 18.3 17.4 13.7 5.3 2.8
FWSVD Model Size (billion) 6.74 4.79 2.95 1.54 0.96 0.47
GSM8K Acc (%) 66.4 62.7 62.5 56.5 1.5 1.9
MATH Acc (%) 20.6 19.2 17.6 14.2 1.8 1.5
Basel Model Size (billion) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.4 63.8 61.6 56.2 50.2 40.2 34.2
MATH Acc (%) 20.6 19.7 17.6 15.5 13.4 10.0 7.8
Таблица 5: Pass@1 accuracy and model size of Llama 2-7B compressed with various pruning algorithms on the mathematical reasoning task.
FLAP Model Size (billion) 6.74 5.39 4.72 4.04 3.37 2.70 1.35
GSM8K Acc (%) 66.4 55.0 40.7 24.8 9.10 0 0
MATH Acc (%) 20.6 14.5 8.4 4.2 1.5 0 0
Wanda Model Size (billion) 6.74 5.39 4.72 4.04 3.37 2.70 1.35
GSM8K Acc (%) 66.4 61.0 44.6 19.0 0.5 0.3 0
MATH Acc (%) 20.6 16.7 12.1 3.5 0.1 0.1 0
Таблица 6: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms on the mathematical reasoning task.
SVD Model Size (billion) 13.02 9.70 6.10 3.27 2.07 1.01
GSM8K Acc (%) 72.7 69.5 63.5 50.0 26.9 6.7
MATH Acc (%) 22.2 20.8 17.8 10.8 5.2 2.2
FWSVD Model Size (billion) 13.02 9.24 5.67 2.93 1.79 0.83
GSM8K Acc (%) 72.7 67.9 63.9 51.9 2.4 3.9
MATH Acc (%) 22.2 20.3 18.0 12.4 1.2 1.9
Basel Model Size (billion) 13.02 9.75 6.13 3.32 2.14 1.12 0.68
GSM8K Acc (%) 72.7 67.9 64.4 55.0 50.3 41.8 37.9
MATH Acc (%) 22.2 20.9 18.7 15.5 13.1 10.4 8.0
Таблица 7: Pass@1 accuracy and model size of Llama 2-13B compressed with various pruning algorithms on the mathematical reasoning task.
FLAP Model Size (billion) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
GSM8K Acc (%) 72.7 60.7 51.2 4.6 0.2 0 0
MATH Acc (%) 22.2 14.5 10.2 4.5 0.4 0 0
Wanda Model Size (billion) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
GSM8K Acc (%) 72.7 66.0 50.3 0.1 0 0 0
MATH Acc (%) 22.2 17.9 12.0 0.4 0.1 0 0
Таблица 8: Pass@1 accuracy and model size of Llama 2-7B compressed with various low-rank algorithms on the code generation task.
SVD Model Size (billion) 6.74 5.02 3.18 1.73 1.11 0.56
HumanEval Acc (%) 23.8 20.7 20.1 9.1 4.9 3.7
MBPP Acc (%) 27.4 21.8 18.6 9.6 2.0 0.4
FWSVD Model Size (billion) 6.74 4.84 3.01 1.58 0.99 0.49
HumanEval Acc (%) 23.8 22.0 20.1 11.6 4.9 0
MBPP Acc (%) 27.4 24.4 17.4 10.4 0 0.6
Basel Model Size (billion) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
HumanEval Acc (%) 23.8 22.0 20.7 14.6 12.8 7.9 7.3
MBPP Acc (%) 27.4 26.6 18.6 12.2 8.8 7.4 4.0
Таблица 9: Pass@1 accuracy and model size of Llama 2-7B compressed with various pruning algorithms on the code generation task.
FLAP Model Size (billion) 6.74 5.40 4.72 4.04 3.37 2.70 1.35
HumanEval Acc (%) 23.8 17.7 9.1 4.3 1.2 0.6 0
MBPP Acc (%) 27.4 22.0 14.4 2.8 1.4 0.2 0
Wanda Model Size (billion) 6.74 5.40 4.72 4.04 3.37 2.70 1.35
HumanEval Acc (%) 23.8 18.9 15.9 3.7 0 0 0
MBPP Acc (%) 27.4 24.4 19.4 7.2 0 0 0
Таблица 10: Pass@1 accuracy and model size of Llama 2-13B compressed with various low-rank algorithms on the code generation task.
SVD Model Size (billion) 13.02 9.70 6.10 3.27 2.07 1.01
HumanEval Acc (%) 27.4 18.9 18.3 3.0 3.7 0.6
MBPP Acc (%) 30.0 25.4 18.2 10.6 1.4 0.6
FWSVD Model Size (billion) 13.02 9.31 5.73 2.97 1.83 0.85
HumanEval Acc (%) 27.4 26.2 20.1 8.5 3.7 0
MBPP Acc (%) 30.0 27.2 21.6 12.2 0.8 0
Basel Model Size (billion) 13.02 9.75 6.13 3.32 2.14 1.12 0.68
HumanEval Acc (%) 27.4 26.2 22.0 15.2 12.8 7.9 7.3
MBPP Acc (%) 30.0 27.8 20.6 13.6 10.8 6.4 3.6
Таблица 11: Pass@1 accuracy and model size of Llama 2-13B compressed with various pruning algorithms on the code generation task.
FLAP Model Size (billion) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
HumanEval Acc (%) 27.4 17.1 15.9 5.5 2.4 0.6 0
MBPP Acc (%) 30.0 20.8 16.0 5.6 2.8 0.6 0
Wanda Model Size (billion) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
HumanEval Acc (%) 27.4 16.5 12.2 0 0 0 0
MBPP Acc (%) 30.0 21.8 13.8 0 0 0 0
Таблица 12: Perplexity and model size of Llama-7B compressed with various low-rank algorithms on WikiText-2. The results of SVD, FWSVD, and SVD-LLM are cited from (svdllm).
SVD Model Size (billion) 6.74 5.10 3.88 2.68 1.29
Perplexity 5.68 20061 52489 105474 687291
FWSVD Model Size (billion) 6.74 5.10 3.88 2.68 1.29
Perplexity 5.68 1727 18156 32194 96872
SVD-LLM Model Size (billion) 6.74 5.10 3.88 2.68 1.29
Perplexity 5.68 7.73 9.27 15.00 31.79
Basel Model Size (billion) 6.74 5.15 4.08 3.21 1.83 1.22 0.67
Perplexity 5.68 6.11 6.68 6.99 7.93 9.21 10.45
Таблица 13: Pass@1 accuracy and model size of Llama 2-7B compressed and 8-bit-quantized by various algorithms for mathematical reasoning.
8-bit FLAP Model Size (GB) 6.74 5.39 4.72 4.04 3.37 2.70 1.35
GSM8K Acc (%) 66.0 53.4 39.1 22.7 7.4 0 0
MATH Acc (%) 20.3 13.5 7.2 4.4 1.0 0 0
8-bit Wanda Model Size (GB) 6.74 5.39 4.72 4.04 3.37 2.70 1.35
GSM8K Acc (%) 66.0 59.7 44.4 15.8 0.3 0.2 0
MATH Acc (%) 20.3 16.3 11.3 2.7 0.1 0 0
8-bit Basel Model Size (GB) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.0 62.1 59.6 54.1 49.7 41.1 35.3
MATH Acc (%) 20.3 18.2 18.1 15.7 13.1 10.1 7.5
Таблица 14: Pass@1 accuracy and model size of Llama 2-13B compressed and 8-bit-quantized by various algorithms for mathematical reasoning.
8-bit FLAP Model Size (GB) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
GSM8K Acc (%) 72.7 59.5 53.9 6.1 0.2 0 0
MATH Acc (%) 21.8 13.9 9.4 4.8 0.1 0 0
8-bit Wanda Model Size (GB) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
GSM8K Acc (%) 72.7 66.4 50.5 0.2 0 0 0
MATH Acc (%) 21.8 17.0 10.7 0.5 0.1 0 0
8-bit Basel Model Size (GB) 13.02 9.75 6.13 3.32 2.14 1.12 0.68
GSM8K Acc (%) 72.7 67.4 61.9 53.4 51.3 40.7 35.9
MATH Acc (%) 21.8 19.3 17.5 14.2 12.5 9.6 7.7
Таблица 15: Pass@1 accuracy and model size of Llama 2-7B compressed and 8-bit-quantized by various algorithms for code generation.
8-bit FLAP Model Size (GB) 6.74 5.39 4.72 4.04 3.37 2.70 1.35
HumanEval Acc (%) 23.8 16.5 9.1 3.7 1.2 0 0
MBPP Acc (%) 26.2 20.8 12.4 3.4 1.2 0.2 0
8-bit Wanda Model Size (GB) 6.74 5.39 4.72 4.04 3.37 2.70 1.35
HumanEval Acc (%) 23.8 16.5 13.4 3.7 0 0 0
MBPP Acc (%) 26.2 22.2 18.2 4.6 0 0 0
8-bit Basel Model Size (GB) 6.74 5.14 3.28 1.21 0.67 0.43
HumanEval Acc (%) 23.8 20.1 19.5 11.6 9.1 7.3
MBPP Acc (%) 26.2 22.4 17.2 8.2 4.6 3.2
Таблица 16: Pass@1 accuracy and model size of Llama 2-13B compressed and 8-bit-quantized by various algorithms for code generation.
8-bit FLAP Model Size (GB) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
HumanEval Acc (%) 25.0 17.7 7.9 4.9 3.2 0.6 0
MBPP Acc (%) 29.0 18.8 12.0 6.8 3.2 0.4 0
8-bit Wanda Model Size (GB) 13.02 10.41 9.11 7.81 6.51 5.21 3.90
HumanEval Acc (%) 25.0 16.5 12.2 0 0 0 0
MBPP Acc (%) 29.0 20.6 14.2 0 0 0 0
8-bit Basel Model Size (GB) 13.02 9.75 6.13 3.32 2.14 1.12 0.68
HumanEval Acc (%) 25.0 25.0 15.9 12.2 11.0 9.1 6.1
MBPP Acc (%) 29.0 26.4 18.4 10.4 9.2 5.4 3.4
Таблица 17: Pass@1 accuracy and model size of Llama 2-7B and -13B 4-bit-quantized by QLoRA for mathematical reasoning and code generation.
Model Task Model size (GB) Accuracy (%)
4-bit QLoRA 7B GSM8K 3.45 54.1
MATH 3.45 12.6
HumanEval 3.45 12.8
MBPP 3.45 16.2
4-bit QLoRA 13B GSM8K 6.63 58.8
MATH 6.63 13.2
HumanEval 6.63 13.4
MBPP 6.63 18.6
Таблица 18: Pass@1 accuracy and model size of 4-bit-quantized Llama 2-7B and Llama 2-13B for mathematical reasoning and code generation.
Model Task Model size (GB) Accuracy (%)
4-bit-quantized 7B GSM8K 3.37 39.2
MATH 3.37 8.8
HumanEval 3.37 6.7
MBPP 3.37 10.8
4-bit-quantized 13B GSM8K 6.51 52.8
MATH 6.51 12.5
HumanEval 6.51 8.5
MBPP 6.51 14.2
Таблица 19: Ablation study: Impact of varying the additional dimension in Basel on the compression of Llama 2-7B for the mathematical reasoning task.
Addt’l Dim 32 Model Size (billion) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.4 63.8 61.6 56.2 50.2 40.2 34.2
MATH Acc (%) 20.6 19.7 17.6 15.5 13.4 10.0 7.8
Addt’l Dim 0 Model Size (billion) 6.74 5.06 3.20 1.75 1.13 0.59
GSM8K Acc (%) 66.4 64.4 61.6 55.4 48.3 34.9
MATH Acc (%) 20.6 19.3 17.4 15.1 12.2 7.1
Таблица 20: Ablation study: Impact of varying the pruning times in Basel on the compression of Llama 2-7B for the mathematical reasoning task.
100 times Model Size (billion) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.4 63.8 61.6 56.2 50.2 40.2 34.2
MATH Acc (%) 20.6 19.7 17.6 15.5 13.4 10.0 7.8
2 times Model Size (billion) 6.74 5.11 3.27 1.82 1.20 0.65
GSM8K Acc (%) 66.4 62.9 59.1 54.4 47.2 32.1
MATH Acc (%) 20.6 18.6 18.1 14.5 11.0 6.2
Таблица 21: Ablation study: Pass@1 accuracy and model size of Llama 2-7B Basel (proposed) and Basel (free basis) on the mathematical reasoning task.
Basel (Proposed) Model Size (billion) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.4 63.8 61.6 56.2 50.2 40.2 34.2
MATH Acc (%) 20.6 19.7 17.6 15.5 13.4 10.0 7.8
Basel (Free Basis) Model Size (billion) 6.74 5.14 4.14 2.51 1.21 0.67 0.43
GSM8K Acc (%) 66.4 61.6 59.8 56.7 44.7 29.0 16.8
MATH Acc (%) 20.6 17.9 17.1 14.9 10.3 6.3 2.6
Таблица 22: Ablation study: Impact of incorporating $L1$ regularization in Basel on the compression of Llama 2-7B for the mathematical reasoning task.
Basel (Proposed) Model Size (billion) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.4 63.8 61.6 56.2 50.2 40.2 34.2
MATH Acc (%) 20.6 19.7 17.6 15.5 13.4 10.0 7.8
Basel ($L1$ norm, $\lambda=0.02$) Model Size (billion) 6.74 5.14 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.4 61.8 61.2 55.2 49.4 39.3 33.7
MATH Acc (%) 20.6 18.8 17.8 15.3 12.5 8.9 8.1
Basel ($L1$ norm, $\lambda=0.1$) Model Size (billion) 6.74 5.13 3.28 1.83 1.21 0.67 0.43
GSM8K Acc (%) 66.4 63.2 61.9 54.1 50.0 40.9 21.2
MATH Acc (%) 20.6 19.5 17.5 15.0 12.4 8.8 3.8